OpenStudy (anonymous):

1280u^10y^7/100u^3y^5..... simplify by removing factors of 1

OpenStudy (anonymous):

\[\frac{1280u^{10}y^7}{100u^3y^5}=12.8u^{10-3}y^{7-5}\]

OpenStudy (anonymous):

or \[12.8u^7y^2\]

OpenStudy (anonymous):

it says its wrong?

OpenStudy (anonymous):

maybe it didn't like the 12.8 try \[\frac{64u^7y^2}{5}\]

OpenStudy (anonymous):

64u^7y^2/5 like that write it?

hero (hero):

I posted a link

OpenStudy (anonymous):

got it